OpenStudy (anonymous):

Determine the number of solutions for the quadratic function f(x) = -x2 + 3x - 9?

OpenStudy (anonymous):

I said there is none

OpenStudy (anonymous):

If by solution you mean roots, then there are 2 imaginary roots.

OpenStudy (anonymous):

how is there 2 distinct real solutions...?

OpenStudy (raden):

just check the discriminant value (D), with D = b^2 - 4ac if D > 0 then there are 2 distinc real solutions if D = 0 then there is one real solution (the roots are same) if D < 0 then there are 2 imaginary solutions (the roots not real)

OpenStudy (anonymous):

So it is 2 distinct real solutions! Thanks guys!

OpenStudy (anonymous):

no

OpenStudy (raden):

i didnt say that --"

OpenStudy (anonymous):

1 solution with a double root?

OpenStudy (anonymous):

if D < 0 then there are 2 imaginary solutions (the roots not real)

OpenStudy (raden):

f(x) = -x2 + 3x - 9 according information above, a = -1, b = 3, c = -9 D = b^2 - 4ac D = 3^2 - 4(-1)(-9) D = 9 - 36 D = -27 that means D < 0 take your conclution, now

OpenStudy (anonymous):

I'm confused as pellet right now. No real solutions or 1 solution (double root) or 2 distinct real solutions or 3 solutions

OpenStudy (anonymous):

Those are all my guesses :\

OpenStudy (anonymous):

No real solution.

OpenStudy (anonymous):

no solution is your case

OpenStudy (anonymous):

I was right in the first place. What the...